OpenStudy (anonymous):

What do you get when you simplify 168/280

OpenStudy (anonymous):

0.6

OpenStudy (anonymous):

I multiplied 7/8 by 24/35 and I got 168/280

OpenStudy (anonymous):

You can tell from the last two digits that both are divisible by 4. (168/4) / (280/4) = 42/70 We can divide these both by 2. (42/2)/(70/2) = 21/35

OpenStudy (anonymous):

So you can get 0.6 when simplified

OpenStudy (anonymous):

Do you have multiple choices?

OpenStudy (anonymous):

No

OpenStudy (anonymous):

Oh, I thought you wanted a simplified fraction. Either way, dividing gives you 0.6, so yes.

OpenStudy (anonymous):

My answer was 24/35 because 3/5 times 7/8 is 24/35 and when u check you multiply that by 7/8 and you simplify your answer to 3/5

OpenStudy (anonymous):

It's 0.6, yes :)